OpenStudy (anonymous):

Is ΔABC with vertices A(–2, 3), B(2, 0), and C(–1, –4) a right triangle? ΔABC is a right triangle because . ΔABC is a right triangle because . ΔABC is a right triangle because . ΔABC is not a right triangle.

OpenStudy (anonymous):

Someone please help me? I don't know how to do this problem!

OpenStudy (anonymous):

use a^2=b^2=c^2

OpenStudy (anonymous):

I honestly still don't get it. Math isn't my specialty.

OpenStudy (anonymous):

@nonamenever

OpenStudy (anonymous):

well first find the lenghts of the tri then use this equation if its true then its a tri